An electron in the ground state of a silicon atom has the quantum numbers (3, 1, -1, +1⁄2) and absorbs infrared radiation of wavelength 1095 nm.

(i) To which energy level does the electron move?

(ii) Calculate the frequency of this radiation.


1
Expert's answer
2021-01-13T13:52:47-0500

i)

"v" = R ( "\frac{1}{n_1^2}" - "\frac{1}{n_2^2}" ) , "v" = frequency ,R = 109700 , wavelength =1095nm = 1095×10-7

"v" = 1/λ

1/λ = R("\frac{1}{n_1^2}" - "\frac{1}{n_2^2}" )


"\frac{1}{1095×10^-⁷}" = 109700 ( "\frac{1}{3²}" - "\frac{1}{n_2^2}" )


"\frac{1}{1095×10^-⁷}" = 109700 ("\frac{n_2^2-9}{9n_2^2}" )


"\frac{1}{1095×10^-⁷×109700}" = "\frac{n_2^2-9}{9n_2^2}"


"\frac{1}{12.01215}" = "\frac{n_2^2-9}{9n_2^2}"


9n"_2^2" = 12.01215(n"_2^2" - 9)


9n"_2^2" = 12.01215n"_2^2" - 108.10935


12.01215n"_2^2" - 9n"_2^2" = 108.10935


3.01215 n"_2^2" = 108.10935


n"_2^2" = 108.10935/3.01215


n"_2^2" = 36


n2 = 6


ii) "v" = c/λ

"v" = (3×10⁸) / (1095×10-9)

"v" = ( 3× 1017 ) / 1095

"v" = 2.73 × 1014